Stratford Company distributes a lightweight lawn chair that sells for $60 per unit. Variable expenses are $24 per unit, and fixed expenses total $777,600 annually.

Required:

Answer the following independent questions:

1.

What is the product’s CM ratio?

    

2.

Use the CM ratio to determine the break-even point in sales dollars.

      

3.

The company estimates that sales will increase by $85,000 during the coming year due to increased demand. By how much should net operating income increase?

    

4.

Assume that the operating results for last year were as follows:

  Sales $ 2,160,000
  Less: Variable expenses 864,000
  Contribution margin 1,296,000
  Less: Fixed expenses 777,600
  Net operating income $ 518,400
a.

Compute the degree of operating leverage at the current level of sales. (Round your answer to 1 decimal place.)

           

b.

The president expects sales to increase by 60% next year. By how much should net operating income increase?

             

5-a.

Refer to the original data. Assume that the company sold 38,000 units last year. The sales manager is convinced that a 10% reduction in the selling price, combined with a $144,000 increase in advertising expenditures, would increase annual unit sales by 30%. Prepare two contribution format income statements: one showing the results of last year’s operations, and one showing what the results of operations would be if these changes were made. (Do not round intermediate calculations. Round "Per Unit" answers to 2 decimal places.)

      

5-b. Would you recommend that the company do as the sales manager suggests?
Yes
No
6.

Refer to the original data. Assume again that the company sold 38,000 units last year. The president feels that it would be unwise to change the selling price. Instead, he wants to increase the sales commission by $2 per unit. He thinks that this move, combined with some increase in advertising, would double annual unit sales. By how much could advertising be increased with profits remaining unchanged?
Answer #1

Contribution per unit = Sales Value - Variable cost

60 - 24 = 36

(1) Contribution Margin Ratio = Contribution*100/Sales

36*100/60 = 60%

(2) Break even Sales in Dollars = Fixed Cost/ Contribution Margin Ratio

777600/60% = 1296000

(3) Increase in net operating income because of increase in Sales-

85000*60% = 51000

(4)(a) Degree of Operating Leverage = Contribution/profit

1296000/518400 = 2.5

(b) Contribution Margin Ratio= 1296000*100/2160000 = 60%

Increase in sales if it increase by 60% = 2160000*60% = 1296000

increase in net operating profit is sales increase by 60%-

1296000*60% = 777600.

(5)(a)

Last Year Next Year
Units 38000 49400(38000+30%)
Sales 2280000(38000*60) 2667600(49400*54)
Variable Cost 912000 1185600
Contribution Margin 1368000 1482000
Fixed Cost 777600 777600
Advertisement 144000
Net Operating Income 590400 560400
Result Decreased by 30000

(5)(b) No

(6)

Before Change After Change
Units 38000 76000
Sales 2280000(38000*60) 4560000(76000*60)
Variable Cost 912000 1976000(76000*26)
Contribution Margin 1368000 2584000
Fixed Cost 777600 777600
Advertisement x
Net Operating Income 590400 590400
Result Decreased by 30000

2584000-(777600+x) = 590400

x = 1216000.